# Deep Learning-Driven Malware Classification with API Call Sequence Analysis and Concept Drift Handling

### v2

@@ Abstract
Malware classification in dynamic environments presents a significant challenge due to concept drift, where the statistical properties of malware data evolve over time, complicating detection efforts. To address this issue, we propose a deep learning framework enhanced with a genetic algorithm to improve malware classification accuracy and adaptability. Our approach incorporates mutation operations and fitness score evaluations within genetic algorithms to continuously refine the deep learning model, ensuring robustness against evolving malware threats. Experimental results demonstrate that this hybrid method significantly enhances classification performance and adaptability, outperforming traditional static models. Our proposed approach offers a promising solution for real-time malware classification in ever-changing cybersecurity landscapes.
@@ 1 Introduction
Our proposed model leverages the power of neural networks—specifically Artificial Neural Networks (ANNs), Convolutional Neural Networks (CNNs), and Recurrent Neural Networks (RNNs)—to classify malware samples into distinct families based on their behavioral signatures, extracted primarily through $n$-gram analysis of API call sequences. These sequences, derived from dynamic execution in sandbox environments, encapsulate the runtime interactions of PE files with the operating system, offering a rich feature set for discerning malicious intent. However, the true innovation of our approach lies in its integration of GAs to address concept drift. Inspired by natural selection, GAs enable continuous model evolution by applying mutation operations and fitness-based selection to adapt feature sets and optimize classification performance in real time vivekanandan2011mining . This dual strategy ensures that our system not only excels in categorizing known malware but also dynamically adjusts to emerging threats, a capability critical for safeguarding systems in today’s volatile cyber landscape.
The motivation behind this research stems from the escalating arms race between malware developers and cybersecurity defenders. Static models, while computationally efficient, lack the flexibility to counter the rapid evolution of malware, as evidenced by studies highlighting performance degradation over time chen2023overkill . Conversely, purely DL-based approaches, though powerful in pattern recognition, often struggle with generalization across drifting data distributions without retraining jameel2020critical . By marrying DL’s robust feature extraction with GAs’ evolutionary adaptability, we aim to bridge this gap, offering a proactive, scalable solution that anticipates and responds to shifts in malware characteristics. This work builds on prior efforts to handle concept drift garcia2023effectiveness ; fernando2024fesad while pushing the boundaries of malware classification through a focus on API call sequences—a dynamic, behavior-driven indicator of malicious activity.
@@
The remainder of this paper is structured as follows: Section 2 elucidates the foundational concepts of concept drift, its variants, and the pivotal role of GAs and sandbox environments in addressing it. Section 3 reviews related work, situating our contribution within the broader research landscape. Section 4 delineates the data collection and preprocessing pipeline critical to our experiments. Section 5 presents our integrated DL and GA-based framework, detailing its phases from preprocessing to concept drift handling. Section 6 describes the experimental setup—including API key acquisition, data extraction, and model training—and provides an in-depth analysis of results across diverse DL algorithms paired with GAs. Section 7 benchmarks our approach against state-of-the-art techniques, highlighting its strengths and novelty. Section 8 discusses threats to validity, ensuring a rigorous evaluation of our findings. Finally, Section 9 summarizes our contributions and outlines avenues for future research, reinforcing the transformative potential of our methodology in fortifying cybersecurity defenses.
@@ 2 Foundational Concepts for Malware Analysis
Concept Drift: Definition and Taxonomy
@@
Detecting and adapting to concept drift is paramount in dynamic environments like malware detection, where failure to adapt leads to degraded performance (e.g., increased false negatives). Techniques range from passive monitoring of prediction errors to proactive model updates, as detailed next.
Techniques to Handle Concept Drift
Maintaining model efficacy amidst concept drift requires a suite of adaptive strategies tailored to the drift’s characteristics. Below, we outline key techniques, emphasizing their applicability to malware classification:
@@
Each strategy offers trade-offs: online learning excels in real-time settings but may overfit noise, while ensembles provide stability at higher computational cost. The choice depends on drift type, data volume, and application constraints, with our study favoring GAs for their evolutionary adaptability.
Genetic Algorithms for Concept Drift
Genetic algorithms (GAs), inspired by Darwinian evolution, are optimization heuristics that emulate natural selection through mutation, crossover, and selection vivekanandan2011mining . In machine learning, GAs address concept drift by evolving solutions to align with changing environments. Operating on a population of candidate solutions (e.g., feature sets, models), GAs iterate as follows: 1. Initialization: Generate a random population. 2. Fitness Evaluation: Score individuals using a fitness function (e.g., classification accuracy). 3. Selection: Favor high-fitness individuals (e.g., via tournament selection). 4. Crossover: Combine parent solutions to produce offspring. 5. Mutation: Introduce random changes to maintain diversity. 6. Replacement: Update the population with new generations.
For concept drift, GAs adapt models dynamically:
@@
    Feature Selection: GAs identify relevant features (e.g., API $n$-grams) for the current drift state, evolving a chromosome of binary feature masks (1 = selected, 0 = discarded). Fitness might be F1-score on recent data.
@@ 2.1 Sandboxing in Malware Analysis
A prominent implementation of these principles is Cuckoo Sandbox, an open-source, automated malware analysis system accessible at https://github.com/cuckoosandbox/ cuckoo_repo . Cuckoo Sandbox exemplifies the sandbox paradigm by providing a robust framework for analyzing suspicious files and URLs within a virtualized setting. It leverages virtualization technologies (e.g., VirtualBox, VMware, or KVM) to instantiate isolated guest environments—typically Windows instances—where PE files are executed safely. Cuckoo’s architecture comprises a host system that orchestrates analysis and a guest virtual machine (VM) where the malware is detonated. During execution, it captures detailed telemetry, including API call sequences, network traffic (via packet capture), and system modifications, which are compiled into comprehensive JSON reports for subsequent investigation jamalpur2018dynamic . These reports are instrumental in our study, as they provide the $n$-gram API call data driving our deep learning-based classification.
To deploy Cuckoo Sandbox effectively, specific system requirements must be met. The host operating system should be a Linux distribution (e.g., Ubuntu 18.04 LTS, Debian, or CentOS) due to its stability and compatibility with virtualization tools. Hardware demands include a multi-core CPU (e.g., Intel i7 or equivalent) to support concurrent VM execution, ample RAM (minimum 8 GB, recommended 16 GB or more) to accommodate multiple guests, and substantial disk space (at least 500 GB, ideally SSD-backed) to store malware samples, VM images, and analysis logs. Virtualization software such as VirtualBox, VMware Workstation, or KVM is mandatory to create the isolated execution environments, with configuration options to emulate diverse Windows versions (e.g., Windows 10, 7) matching target malware’s operational context. Additionally, Python 3.x is required, along with dependencies like libvirt and volatility, to support Cuckoo’s scripting and memory forensics capabilities. Proper network configuration—including a virtual NIC bridged to a monitored subnet—ensures safe capture of network-based behaviors without risking external exposure.
@@ 4 Data Collection and Preprocessing
In this section, we discuss the process of data collection from VirusShare virusshare and VirusTotal virustotal .
@@ 6 Implementation and Results
In this section, we discuss the experiment setup and analysis of the contained result after the experiment.
@@ 6.1 Experimental Setup
Our experimental setup aimed at evaluating the effectiveness of machine-learning techniques in malware classification. It consists of the following components:
@@ 8 Benign 1724 6910 8634
For the concept drift handling using genetic algorithms, we used a total of 22054 samples, including 4410 test samples and 17644 train samples, across various malware types as well as Benign samples as shown in Table 5 as dataset (Dataset 1). This dataset was collected from VirusShare from January to June 2023 release.
@@
To summarize, we can say that the CNN model outperforms the ANN and RNN models in terms of achieving lower loss and higher accuracy. The ANN model also shows promising results, with a steady improvement in performance with more epochs. However, the RNN model struggles, showing minimal improvement in loss and accuracy throughout training. These results suggest that CNN model is most suitable for this task, followed by ANN model, while RNN model may not be well-suited for this problem, i.e., classifying the malwares with concept drift.
@@ 8 Benign 1720 6880 8600
In contrast, Table 11 presents the performance of these models on an older dataset from VirusShare, released between January and May 2020, highlighting the increase in validation accuracy as well as the decrease in validation loss. Moreover, Table 12 demonstrates the impact of concept drift on the same old dataset (Dataset 2), showing how the models perform when trained on data that has evolved since its original release. These findings suggest that utilising concept drift in the malware dataset can lead to improved testing accuracy and reduced loss, indicating the effectiveness of the concept drift malware classifier.
@@ 7 Comparison with Related work
Table 13: Quantitative comparison of some deep learning-based malware detection and classification techniques with concept drift
@@
2 Fernando et al. fernando2024fesad Genetic Algorithms Elderan dataset 2720 API Calls
@@ 3 Proposed Work Genetic Algorithms VirusShare 22054 API Calls
Fernando and Komninos fernando2024fesad introduced the feSAD ransomware detection framework, which leverages machine learning to adapt to concept drift. The framework aims at enhancing ransomware detection rates by calibrating drift thresholds and identifying abnormal drift samples. Compatible with various machine learning algorithms, it has demonstrated effectiveness in detecting ransomware amidst concept drift and can be tailored for different malware types. Experimental results, including 720 ransomware samples and 2000 benign samples from the Elderan dataset, show high detection rates and precision, especially with the Random Forest algorithm. Komninos highlighted the Random Forest’s stability in ransomware detection. However, they noted that in test batch 2, a high volume of ransomware samples caused abnormal concept drift, leading to reduced detection rates and statistical drift, suggesting system retraining to avoid detection decline.
@@ 10 Code and Dataset Availability
The research code is available on GitHub at the following link: https://github.com/bishwajitprasadgond/MalClassCD. For access to the dataset used in this research, please send a request via email to bishwajitprasadgond@gmail.com.

### v4

@@ Abstract
Malware classification in dynamic environments presents a significant challenge due to concept drift, where the statistical properties of malware data evolve over time, complicating the detection effort. To address this issue, we propose a deep learning framework enhanced with a genetic algorithm to improve malware classification accuracy and adaptability with concept drift handeling. Our approach incorporates mutation operations and fitness score evaluations within genetic algorithms to continuously refine the deep learning model, ensuring robustness against evolving malware threats. Experimental results demonstrate that this hybrid method significantly enhances classification performance and adaptability, outperforming traditional static models. Our proposed approach offers a promising solution for real-time malware classification in ever-changing cybersecurity landscapes.
@@ 1 Introduction
Our proposed model leverages the power of neural networks, specifically Artificial Neural Networks (ANN), Convolutional Neural Networks (CNN) and Recurrent Neural Networks (RNNs) to classify malware samples into distinct families based on their behavioral signatures, extracted primarily through $n$ gram analysis of API call sequences. These sequences, derived from dynamic execution in sandbox environments, encapsulate the runtime interactions of PE files with the operating system, offering a rich feature set for discerning malicious intent. However, the true innovation of our approach lies in its integration of GAs to address concept drift. Inspired by natural selection, GAs enable continuous model evolution by applying mutation operations and fitness-based selection to adapt feature sets and optimize classification performance in real time vivekanandan2011mining . This dual strategy ensures that our system not only excels in categorizing known malware but also dynamically adjusts to emerging threats, a capability critical for safeguarding systems in today’s volatile cyber landscape.
The motivation behind this research stems from the escalating arms race between malware developers and cybersecurity defenders. Static models, while computationally efficient, lack the flexibility to counter the rapid evolution of malware, as evidenced by studies highlighting performance degradation over time chen2023overkill . Conversely, purely DL-based approaches, though powerful in pattern recognition, often struggle with generalization across drifting data distributions without retraining jameel2020critical . By marrying DL’s robust feature extraction with GAs’ evolutionary adaptability, we aim at bridging this gap, offering a proactive, scalable solution that anticipates and responds to shifts in malware characteristics. This work builds on prior efforts to handle concept drift garcia2023effectiveness ; fernando2024fesad while pushing the boundaries of malware classification through a focus on API call sequences—a dynamic, behavior-driven indicator of malicious activity.
@@
The remainder of this paper is structured as follows: Section 2 elucidates the foundational concepts of concept drift, its variants, and the pivotal role of GAs and sandbox environments in addressing it. Section 3 reviews the related work, situating our contribution within the broader research landscape. Section 4 delineates the data collection and preprocessing pipeline critical to our experiments. Section 5 presents our integrated DL and GA-based framework, detailing its phases from preprocessing to concept drift handling. Section 6 describes the experimental setup—including API key acquisition, data extraction, and model training—and provides an in-depth analysis of results across diverse DL algorithms paired with GAs. Section 7 benchmarks our approach against state-of-the-art techniques, highlighting its strengths and novelty. Section 8 discusses threats to validity, ensuring a rigorous evaluation of our findings. Finally, Section 9 summarizes our contributions and outlines avenues for future research, reinforcing the transformative potential of our methodology in fortifying cybersecurity defenses.
@@ 2 Foundational Concepts for Malware Analysis
2.1 Concept Drift: Definition and Taxonomy
@@
Detecting and adapting to concept drift is paramount are dynamic environments like malware detection, where failure to adapt leads to degraded performance (e.g., increased false negatives). The existing techniques range from passive monitoring of prediction errors to proactive model updates, as detailed in the following subsection.
2.2 Techniques to Handle Concept Drift
Maintaining model efficacy amidst concept drift requires a suite of adaptive strategies tailored to the drift’s characteristics. Below, we outline the key techniques, emphasizing their applicability to malware classification:
@@ 2.2 Techniques to Handle Concept Drift
Each strategy offers trade-offs: online learning excels in real-time settings but may overfit noise, while ensembles provide stability at higher computational cost. The choice depends on drift type, data volume, and application constraints, Our study uses GAs for their evolutionary adaptability.
2.3 Genetic Algorithms for Concept Drift
Genetic algorithms (GAs), inspired by Darwinian evolution, are optimization heuristics that emulate natural selection through mutation, crossover, and selection vivekanandan2011mining . In machine learning, GAs address concept drift by evolving solutions to align with changing environments. Operating on a population of candidate solutions (e.g., feature sets, models), GAs iterate as follows:
@@ 2.3 Genetic Algorithms for Concept Drift
    Initialization: Generate a random population.
  2. 2.
    Fitness Evaluation: Score individuals using a fitness function (e.g., classification accuracy).
  3. 3.
    Selection: Favor high-fitness individuals (e.g., via tournament selection).
  4. 4.
    Crossover: Combine parent solutions to produce offspring.
  5. 5.
    Mutation: Introduce random changes to maintain diversity.
  6. 6.
    Replacement: Update the population with new generations.
For handelling concept drift, GAs adapt models dynamically as follows:
  1. 1.
    Feature Selection: GAs identify the relevant features (e.g., API $n$-grams) for the current drift state, evolving a chromosome of binary feature masks (1 = selected, 0 = discarded). Fitness might be the F1-score on recent data.
@@ 2.4 Sandboxing in Malware Analysis
A prominent implementation of these principles is Cuckoo Sandbox, an open-source, automated malware analysis system accessible at https://github.com/cuckoosandbox/ cuckoo_repo . Cuckoo Sandbox exemplifies the sandbox paradigm by providing a robust framework for analyzing suspicious files and URLs within a virtualized setting. It leverages virtualization technologies (e.g., VirtualBox, VMware, or KVM) to instantiate isolated guest environments—typically Windows instances—where PE files are executed safely. Cuckoo’s architecture comprises of a host system that orchestrates analysis and a guest virtual machine (VM) where the malware is detonated. During execution, it captures detailed telemetry, including API call sequences, network traffic (via packet capture), and system modifications, which are compiled into comprehensive JSON reports for subsequent investigation jamalpur2018dynamic . These reports are instrumental in our study, as they provide the $n$-gram API call data driving our deep learning-based classification.
To deploy Cuckoo Sandbox effectively, specific system requirements must be met. The host operating system should be a Linux distribution (e.g., Ubuntu 18.04 LTS, Debian, or CentOS) due to its stability and compatibility with virtualization tools. Hardware demands include a multi-core CPU (e.g., Intel i7 or equivalent) to support concurrent VM execution, ample RAM (minimum 8 GB, recommended 16 GB or more) to accommodate multiple guests, and substantial disk space (at least 500 GB, ideally SSD-backed) to store malware samples, VM images, and analysis logs. Virtualization software such as VirtualBox, VMware Workstation, or KVM is mandatory to create the isolated execution environments, with configuration options to emulate diverse Windows versions (e.g., Windows 10, 7) matching target malware’s operational context. Additionally, Python 3.12 is required, along with dependencies like libvirt and volatility, to support Cuckoo’s scripting and memory forensics capabilities. Proper network configuration, including a virtual NIC bridged to a monitored subnet, ensures safe capture of network-based behaviors without risking external exposure.
@@ 4 Data Collection and Preprocessing
In this section, we discuss the process of data collection from VirusShare virusshare and VirusTotal virustotal , as shown in Figure  1.
@@ 6 Implementation and Results
In this section, we discuss the experimental setup and analysis of the obtained results.
@@ 6.1 Experimental Setup
Our experimental setup aimed at evaluating the effectiveness of deep-learning techniques in malware classification. It consists of the following components:
@@ 8 Benign 1724 6910 8634
For the malware classification experiment, we used a total of 22054 samples, including 4410 test samples and 17644 train samples, across various malware types as well as Benign samples as shown in Table 5 as dataset (Dataset 1). This dataset was collected from VirusShare from January to June 2023 release and we tested our deep-learning model without concept drift.
@@
Figure 7: Epoch vs Loss and Epoch vs Accuracy for RNN
To summarize, we can say that the CNN model outperforms the ANN and RNN models in terms of achieving lower loss and higher accuracy. The ANN model also shows promising results, with a steady improvement in performance with more epochs. However, the RNN model struggles, showing minimal improvement in loss and accuracy throughout training. These results suggest that CNN model is most suitable for this task, followed by ANN model, while RNN model may not be well-suited for this problem, i.e., classifying the malwares without concept drift.
@@ 8 Benign 1720 6880 8600
In contrast, Table 11 presents the performance of these models on an older dataset from VirusShare, released between January and May 2020, highlighting the decrease in validation accuracy as well as the increase in validation loss. Moreover, Table 12 demonstrates the impact of concept drift on the same old dataset (Dataset 2), showing how the models perform when trained on data that has evolved since its original release. These findings suggest that utilising concept drift in the malware dataset can lead to improved testing accuracy and reduced loss, indicating the effectiveness of the concept drift malware classifier.
@@ 7 Comparison with Related work
Fernando and Komninos fernando2024fesad introduced the feSAD ransomware detection framework, which leverages machine learning to adapt to concept drift. The framework aims at enhancing ransomware detection rates by calibrating drift thresholds and identifying abnormal drift samples. Compatible with various machine learning algorithms, it has demonstrated the effectiveness in detecting ransomware amidst concept drift and can be tailored for different malware types. Experimental results, including 720 ransomware samples and 2000 benign samples from the Elderan dataset, show high detection rates and precision, especially with the Random Forest algorithm. Fernando and Komninos fernando2024fesad highlighted the Random Forest’s stability in ransomware detection. However, they noted that in test batch 2, a high volume of ransomware samples caused abnormal concept drift, leading to reduced detection rates and statistical drift, suggesting system retraining to avoid detection decline.
Table 13: Comparison with some related work
@@
2 Fernando and Genetic Algorithms Elderan dataset 2720 API Calls
Komninosfernando2024fesad
@@ 10 Code and Dataset Availability
The code for this work is available on GitHub at the following link: https://github.com/bishwajitprasadgond/MalClassCD. For access to the dataset used in this research, please send a request via email to bishwajitprasadgond@gmail.com.